# Functional-level Uncertainty Quantification for Calibrated Fine-tuning on LLMs

### v2

@@ 3.4 Training Calibration Loss
The FEU model of the epistemic uncertainty enables us to calibrate the parameter mixtures against predictive accuracy for better aligned output distributions. Specifically for the MoE top-k routers, we design the following calibration loss for training:
$\mathcal{L}_{\mathrm{cal}}=\left\|\mathbbm{1}{\{\text{MixLoRA}(x)=y^{*}\}}-\text{FEU}(x)\right\|^{2},$ (13)
@@
where the first term is an indicator function of whether the model prediction matches the ground truth $y^{*}$ given the prompt $x$. We employ a one-hot definition of the ground truth confidence. When the prediction from current mixture model matches the ground truth, the ground truth confidence is $1$. Otherwise, when the predictions do not match, the ground truth confidence is 00.
3.5 Calibrating the Mixture
@@
Our proposed calibration loss $\mathcal{L}_{\mathrm{cal}}$ improves predictive reliability by adaptively balancing expert exploitation and exploration with functional-level epistemic uncertainty. As shown in Figure 1, $\mathcal{L}_{\mathrm{cal}}$ effectively promotes expert exploitation for correct predictions and expert exploration for incorrect predictions by directly calibrating the functional level epistemic uncertainty to align with the predictive correctness.
Ideally, when the $K$ LoRA experts together capture all the functional relationships across the data distribution with cross entropy during fine-tuning, our proposed loss $\mathcal{L}_{cal}$ also finds proper mixture of LoRA experts conditioned on the input $x$ by conditionally promoting expert exploitation and exploration. This allows the model to select correct functional relationships regarding $x$ to generate an output that better matches the data distribution, which grants calibrated uncertainty estimations.
Load balancing is a common regularization to ensure even utilization across experts with the MoE architecture (Fedus et al., 2022). We follow the load balancing loss $\mathcal{L}_{b}$ proposed by Li et al. (2024) and define our loss function as:
$\mathcal{L}=\mathbf{CE}+\gamma\cdot\mathcal{L}_{b}+\beta\cdot\mathcal{L}_{\mathrm{cal}},$ (14)
@@
where $\mathbf{CE}$ represents cross entropy loss, $\gamma$ and $\beta$ are the hyperparameters of two auxiliary terms. Details about $\mathcal{L}_{b}$ can be found in Appendix A.1.
@@ 4 Related Work
4.1 Parameter-Efficient Fine Tuning for LLMs
Large Language Models (Brown et al., 2020; Hoffmann et al., 2022; Touvron et al., 2023b, c) have shown impressive abilities in handling various natural language processing tasks. Building on these advances, instruction fine-tuning (Chung et al., 2022; Iyer et al., 2022; Zheng et al., 2024) has enhanced LLMs’ capacity to comprehend and follow human instructions, forming the core of modern conversational AI systems (Wu et al., 2023b; Achiam et al., 2023b). However, as LLMs increase in size, the fine-tuning process demands much more time and memory.
To address these challenges, several strategies have been proposed, including parameter-efficient fine-tuning (PEFT) (Mangrulkar et al., 2022), model distillation (Liu et al., 2023; Xiao et al., 2023), quantization (Frantar et al., 2022; Xiao et al., 2022b), and pruning (Ma et al., 2023). Among these, LoRA (Hu et al., 2021b), which leverages low-rank matrix decomposition of linear layer weights, is a widely adopted PEFT technique that boosts model performance without adding computational costs during inference. For example, VeRA (Kopiczko et al., 2023) introduces learnable scaling vectors to modify shared pairs of frozen random matrices across layers, while FedPara (Hyeon-Woo et al., 2021) focuses on low-rank Hadamard products for federated learning settings. Tied-Lora (Renduchintala et al., 2023) applies weight tying to further minimize the number of trainable parameters. AdaLoRA (Zhang et al., 2023) uses Singular Value Decomposition (SVD) to prune less important singular values for efficient updates, and DoRA (Liu et al., 2024) separates pre-trained weights into magnitude and direction components, applying LoRA to update the directional component during fine-tuning, thus reducing the number of parameters to be trained.
4.2 Mixture of Experts
The Mixture-of-Experts concept (Jacobs et al., 1991; Wang et al., 2020), introduced as early as 1991, presented a novel supervised learning framework where multiple networks (experts) specialize in handling distinct subsets of training data. Modern MoE variants adapt this by modifying the traditional feed-forward sub-layer within transformer blocks, incorporating sparsely activated LoRA experts, which allows for significant expansion in model width without a proportional increase in computational overhead.
Different MoE architectures have since emerged, distinguished by their expert sampling and routing strategies. For example, LLaVA-MoLE (Chen et al., 2024) improves token routing to domain-specific experts within transformer layers, reducing data conflicts and consistently outperforming standard LoRA baselines. Other MoE-based approaches include MoRAL (Yang et al., 2024a), which focuses on efficiently adapting LLMs to new domains and tasks for lifelong learning, and LoRAMoE (Dou et al., 2024), which incorporates LoRAs via a router network to mitigate the issue of world knowledge forgetting. PESC (Wu et al., 2024b) transforms dense models into sparse ones through an MoE structure, lowering computational costs. MoE-LoRA (Luo et al., 2024) introduces a new parameter-efficient MoE method using Layer-wise Expert Allocation (MoLA) for transformer models, while MoCLE (Gou et al., 2023) activates task-specific model parameters based on instruction clusters. MixLoRA (Li et al., 2024) implements a high-throughput framework for LoRA MoE training and inference process, constructing LoRAs as stochastic experts to reduce computational overhead while expanding model capacity.
Despite the performance improvements these architecture advancements have brought, the overconfidence problem of fine-tuned models is lacking attention (Xiao et al., 2022a; He et al., 2023; Tian et al., 2023; OpenAI, 2023). Enhancing the uncertainty estimation capabilities of these models is fundamental toward more reliable, interpretable and trustworthy applications of LLMs.
4.3 Uncertainty Quantification in LLMs
Uncertainty quantification has garnered substantial attention in various tasks and domains within neural networks (Gal and Ghahramani, 2015, 2016a; Malinin and Gales, 2018; Ovadia et al., 2019; Dusenberry et al., 2020; Malinin et al., 2021; Lin et al., 2022; Kuhn et al., 2023; Lin et al., 2023). This focus extends to LLMs, where the precise quantification of prediction uncertainty has become a critical area of research (Xiao et al., 2022c; Lin et al., 2022; Mielke et al., 2022; Chen and Mueller, 2023; Duan et al., 2023; Huang et al., 2023). LLMs, particularly in generative tasks, pose unique challenges, especially when it comes to measuring the uncertainty of their outputs (Liu et al., 2019; Malinin and Gales, 2021; Kuhn et al., 2023; Lin et al., 2023). The distinction between aleatoric and epistemic uncertainty was recently examined in the context of LLMs (Hou et al., 2023), though this was approached by ensembling model inputs rather than model instances and did not address fine-tuning tasks specifically.
@@ 4.3 Uncertainty Quantification in LLMs
Established uncertainty quantification methods have been studied in conjunction with the LoRA structure. Monte-Carlo dropout (Gal and Ghahramani, 2016b) interprets dropout in neural networks as approximate Bayesian inference in deep Gaussian processes, allowing uncertainty estimates to be obtained from existing LoRA adapters without modifying them. Checkpoint ensemble (Chen et al., 2017) utilizes predictions from multiple LoRA checkpoints saved during a single fine-tuning process to calibrate uncertainty. Deep ensemble (Lakshminarayanan et al., 2017; Wang et al., 2023; Zhai et al., 2023) combines the predictions from multiple LoRA adapters for better uncertainty calibration. Laplace-LoRA (Yang et al., 2024b) applies Bayesian inference via Laplace approximation to the LoRA parameters after fine-tuning, resulting in improved calibration and uncertainty estimates. However, none of these methods provide calibrations regarding epistemic uncertainty, which is crucial to mitigate overconfidence in the fine-tuning stage given the sparse dataset.
@@ 5 Experiments
5.1 Datasets
We include $5$ multiple-choice question answering benchmarks: OpenBookQA (OBQA, Mihaylov et al. (2018)), ARC-Easy (ARC-E) and ARC-Challenge (ARC-C) from AI2 Reasoning Challenge (Clark et al., 2018), BOOLQ (Clark et al., 2019) and ClimateQA, an expert-annotated domain specific benchmark for climate science. We also use computer science, law, medication and engineering subsets from MMLU dataset (Hendrycks et al., 2020) to evaluate performance under distribution shift. We fine-tune on the publicly available training split and test on the validation split from these benchmarks to evaluate model performance. We report the average model performance over $3$ random runs and the standard deviations in the subscript.
Table 1: Performance comparison of different methods fine-tuned with LlaMA2-7B across four common sense reasoning tasks and a domain-specific task. UQ4CT shows substantial ECE improvements while maintaining high accuracy.
Metrics Methods BoolQ ARC-E ARC-C OBQA ClimateQA
ACC $\uparrow$ LoRA $69.5_{1.93}$ $74.8_{1.39}$ $53.8_{0.6}$ $72.1_{0.87}$ $59.9_{2.13}$
MC Drop $66.8_{3.66}$ $76.8_{1.30}$ $50.9_{2.01}$ $74.8_{1.34}$ $58.2_{2.11}$
Ensemble $66.2_{3.7}$ $71.2_{1.0}$ $47.5_{0.57}$ $75.5_{1.4}$ $59.6_{6.9}$
LA $68.7_{1.32}$ $74.6_{2.11}$ $51.4_{0.83}$ $70.8_{1.24}$ $55.2_{3.29}$
MixLoRA $71.5_{1.05}$ $77.7_{2.27}$ $54.3_{1.07}$ $75.5_{2.91}$ $61.6_{1.76}$
UQ4CT $73.5_{0.52}$ $76.6_{1.30}$ $52.8_{1.77}$ $77.3_{1.36}$ $63.3_{1.74}$
ECE $\downarrow$ LoRA $11.9_{0.78}$ $11.9_{2.04}$ $19.4_{4.75}$ $10.2_{1.07}$ $14.3_{0.64}$
MC Drop $12.2_{0.85}$ $11.9_{1.99}$ $19.8_{4.85}$ $10.9_{0.24}$ $14.3_{0.56}$
Ensemble $7.28_{2.3}$ $9.1_{1.49}$ $10.23_{1.39}$ $8.83_{2.35}$ $13.5_{3.29}$
LA $17.1_{1.72}$ $16.6_{3.7}$ $18.1_{0.5}$ $17.2_{1.2}$ $12.6_{1.9}$
MixLoRA $7.88_{2.09}$ $9.09_{0.81}$ $10.74_{1.07}$ $12.9_{1.99}$ $12.5_{1.32}$
5.2 Experiment Setup
We implement UQ4CT with PyTorch (Paszke et al., 2019), extending the MixLoRA repository in (Li et al., 2024) and compare the average performance in three random runs and report the mean and standard deviation with following baselines. We use the publicly available LLaMA-2-7B-hf (Touvron et al., 2023a) as our base model. In particular, we apply MixLoRA to query, key, value and output layers, together with the feed-fowrward networks in LLaMA-2-7B-hf (gate layer, down layer and up layer). Details are provided in Appendix A.5
@@ 5.2 Experiment Setup
  • Monte Carlo (MC) Dropout (Gal and Ghahramani, 2016b) keeps dropout on at both training and testing time. By performing multiple forward passes, MC dropout randomly shuts down a portion of model nodes, producing ensemble-alike predictions. To combine LoRA fine-tuning with MC dropout, we apply dropout on the input of the LoRA adapter, following the implementation of Mangrulkar et al. (2022).
@@
  • Laplace-LoRA (LA) (Yang et al., 2024b) applies a post-hoc Laplace approximation on fine-tuned LoRA parameters for better uncertainty estimation.
@@
Evaluation Metrics.
We measure the prediction accuracy on the validation set for all $5$ tasks. For uncertainty calibration, we incorporate expected calibration error (ECE, Guo et al. (2017)) with $15$ bins, which measures the alignment between predicted probabilities and empirical accuracy. We also investigate model performance under distribution shift to ensure the model has predictable behavior when given data from other domains as this is a crucial component for real-world applications. Specifically, we test models fine-tuned on OBQA dataset with $4$ domain-specific MMLU subtask ensembles focusing on different professionalities and ARC-E/C datasets to approximate larger and smaller distribution shifts. Metric details are given in Appendix A.6.
5.3 Results
We assess the prediction accuracy and uncertainty calibration of models under both in-distribution and distribution shift scenarios. The in-distribution scenario examines the alignment of the fine-tuned model on the target downstream task, while the distribution shift scenario evaluates the generalizability of the model on novel tasks beyond the fine-tuned domain. These two scenarios combined provides a comprehensive assessment of model robustness in real-world applications where it is essential for the model to excel on its primary task while maintaining the ability to effectively handle unforeseen or out-of-distribution inputs.
@@ 5.3 Results
MixLoRA $75.5_{2.91}$ $58.5_{1.44}$ $69.2_{1.02}$ $35.2_{2.92}$ $30.3_{0.98}$ $35.9_{0.43}$ $40.6_{1.13}$
UQ4CT $77.3_{1.36}$ $58.8_{1.06}$ $65.8_{1.31}$ $36.2_{1.24}$ $34.1_{2.31}$ $35.8_{1.01}$ $40.0_{1.24}$
@@
5.3.1 In-distribution Performance
We first evaluate UQ4CT and baseline models fine-tuned on the $4$ common sense reasoning tasks and the climate question answering task under the in-distribution scenario, where models are trained and evaluated on different splits of the same dataset. Note that one of the key advantages of UQ4CT is that uncertainty calibration happens during the fine-tuning stage with little computational overhead, while other UQ methods require repetitive sampling or other post-hoc complexities.
As shown in Table 1, UQ4CT demonstrates notable improvements in uncertainty calibration across a variety of tasks. Across all benchmarks, UQ4CT maintains competitive accuracy (ACC) compared to the baseline methods. For example, on the BoolQ and ClimateQA datasets, UQ4CT achieves accuracy rates of $73.5\%$ and $63.3\%$, respectively. This empirically demonstrates that UQ4CT is capable of maintaining high accuracy with uncertainty calibration, which assures the gain in UQ performance does not compromise accuracy.
The most substantial performance improvement is observed in the reduction of Expected Calibration Error (ECE). UQ4CT consistently outperforms other methods, reducing ECE by more than $25\%$ on average across the evaluated benchmarks. Unlike other methods where the ECE performance is worsened on the more challenging ARC-C benchmark, UQ4CT achieves an ECE score of $6.1$, showcasing the effectiveness of the calibration.
In addition to experiments on LLaMA-2-7B in the main text, we also present additional experiments on fine-tuning Mistral-7B in Appendix A.2 for more comprehensive evaluation of our method. For both LLaMA-2-7B and Mistral-7B models, UQ4CT consistently shows substantial improvements in uncertainty calibration. The improvements are critical in applications where the model’s confidence must align with its predictive accuracy given limited data, particularly in safety-critical and domain-specific tasks.
5.3.2 Performance Under Distribution Shift
Due to the sparse nature of the fine-tuning data, real world deployment of LLMs often requires the model to be robust to out-of-distribution knowledge (Ouyang et al., 2022; Touvron et al., 2023d, a). Therefore, we evaluate the performance of UQ4CT along with other baseline models fine-tuned on the OBQA dataset under smaller and larger distribution shift scenarios. Similar to the dataset setup in (Yang et al., 2024b), we use ARC-C and ARC-E dataset to simulate smaller distribution shift because the ARC dataset has similar domain focus on general science reasoning, but is generally more challenging and covers a broader range of scientific topics than OBQA. For larger distribution shift, we ensemble the domain-specific MMLU subtasks into $4$ benchmarks focusing on different professionalities: Computer Science (CS), Engineering (Eng), Law and Health. These tasks have very broad coverage of the domain task at various knowledge levels ranging from elementary school to professionals. This domain-specificity demonstrates larger distribution shift from OBQA, which is a general common sense reasoning task. Details of the ensemble is provided in Appendix. A.7.

### v3

@@ 3.3 Calibration Loss
The FLU model provides a principled way to calibrate the mixture parameters against predictive accuracy, enabling better alignment between output distributions and true model confidence. Specifically, for MoE top-$k$ routers, we design the following calibration loss for training:
$\mathcal{L}_{\mathrm{cal}}=\left(\mathbbm{1}{\{\text{MixLoRA}(x)=y^{*}\}}-\text{FLU}(x)\right)^{2}.$ (13)
@@
The first term is an indicator function that equals $1$ if the model prediction matches the ground truth $y^{*}$ for prompt $x$, and 00 otherwise, corresponding to a one-hot definition of ground truth confidence.
This calibration loss directly encourages the functional-level uncertainty (FLU) to reflect the true correctness of the model’s predictions. As shown in Figure 2, when the mixture model predicts correctly, the loss pushes FLU toward $1$ (high confidence); when incorrect, toward 00 (low confidence). In other words, when optimized over the data distribution, the calibration loss pushes FLU to represent the probability of prediction correctness. We formally state this property as follows and present the proof in Appendix A.1:
Proposition 3.2 (Truthfulness of Calibration Loss).
Let the calibration risk be defined as the expectation of the calibration loss over the data distribution:
¯$\bar{\mathcal{L}}_{\mathrm{cal}}=\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}\left(\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}-\mathrm{FLU}(x)\right)^{2}.$
If this calibration risk is optimized over the data distribution $\mathcal{D}$, then the optimal solution is $\mathrm{FLU}(x)=\mathbb{P}(\mathrm{MixLoRA}(x)=y^{*}(x))$; that is, the optimally trained FLU corresponds to the probability that the model’s prediction is correct.
3.4 Fine-tuning with the Total Loss
@@
Our proposed calibration loss $\mathcal{L}_{\mathrm{cal}}$ improves predictive reliability by adaptively balancing expert exploitation and exploration according to functional-level uncertainty. As shown in Figure 2, $\mathcal{L}_{\mathrm{cal}}$ aligns uncertainty with predictive correctness, increasing the router probability of the selected expert for correct predictions (exploitation) and decreasing it for incorrect ones (exploration).
Ideally, when the $K$ LoRA experts collectively capture the relevant functional relationships in the data during fine-tuning via cross-entropy loss, $\mathcal{L}_{\mathrm{cal}}$ further guides the model to select appropriate mixtures of LoRA experts conditioned on the input $\mathbf{x}$. This targeted selection enables the model to match the data distribution more closely and provides more calibrated uncertainty estimates.
To ensure balanced expert utilization, we incorporate a load balancing loss $\mathcal{L}_{b}$ as proposed by Li et al. (2024). Our overall loss function is:
$\mathcal{L}=\mathbf{CE}+\gamma\cdot\mathcal{L}b+\beta\cdot\mathcal{L}_{\mathrm{cal}},$ (14)
@@
where $\mathbf{CE}$ is the cross-entropy loss, and $\gamma$, $\beta$ are hyperparameters for the auxiliary terms. Details of $\mathcal{L}_{b}$ are provided in Appendix A.2.
@@ 4 Related Work
Mixture of LoRA Experts.
Large Language Models (LLMs) have achieved impressive performance across diverse NLP tasks (Brown et al., 2020; Hoffmann et al., 2022; Touvron et al., 2023b, c), with instruction fine-tuning (Chung et al., 2022; Iyer et al., 2022; Zheng et al., 2024) further boosting their adaptability for conversational AI (Wu et al., 2023b; Achiam et al., 2023). However, scaling LLMs increases the resource demands of full fine-tuning. Parameter-efficient fine-tuning (PEFT) methods (Mangrulkar et al., 2022)—such as LoRA (Hu et al., 2021b) and its variants (Kopiczko et al., 2023; Hyeon-Woo et al., 2021; Renduchintala et al., 2023; Zhang et al., 2023; Liu et al., 2024)—reduce adaptation costs by updating a subset of parameters.
Recent advances combine PEFT with the Mixture-of-Experts (MoE) framework (Jacobs et al., 1991; Wang et al., 2020), which sparsely activates expert subnetworks for greater model capacity and specialization. MoE-based LLMs leverage expert routing and parameter-efficient adaptations to target new domains or tasks efficiently. Notably, methods such as MoRAL (Yang et al., 2024b), LoRAMoE (Dou et al., 2024), PESC (Wu et al., 2024b), MoE-LoRA (Luo et al., 2024), and MixLoRA (Li et al., 2024) optimize domain-specific routing, mitigate forgetting, and enable scalable, high-throughput training and inference with mixtures of LoRA experts.
Uncertainty Quantification in LLMs.
@@
Established uncertainty quantification methods have been studied in conjunction with the LoRA structure for LLMs. Monte-Carlo dropout (Gal and Ghahramani, 2016) interprets dropout in neural networks as approximate Bayesian inference in deep Gaussian processes, allowing uncertainty estimates to be obtained from existing LoRA adapters without modifying them. Checkpoint ensemble (Chen et al., 2017) utilizes predictions from multiple LoRA checkpoints saved during a single fine-tuning process to calibrate uncertainty. Deep ensemble (Lakshminarayanan et al., 2017; Wang et al., 2023; Zhai et al., 2023) combines the predictions from multiple LoRA adapters for better uncertainty calibration. Laplace-LoRA (Yang et al., 2024a) applies Bayesian inference via Laplace approximation to the LoRA parameters after fine-tuning, resulting in improved calibration and uncertainty estimates.
Prompt-perturbation and resampling-based approaches have also been explored for uncertainty quantification in LLMs. These methods estimate uncertainty by measuring prediction variability under different prompt formulations or sampled input variants, without altering model parameters (Farquhar et al., 2024; Hou et al., 2023; Gao et al., 2024). This line of work leverages the inherent sensitivity of LLMs to input perturbations as a means to assess model confidence, providing a complementary perspective to parameter-based methods. Ye et al. (2024) benchmark LLMs using conformal prediction, which quantifies uncertainty by constructing prediction sets with guaranteed coverage, where set size directly reflects model uncertainty.
@@ 5 Experiments
Datasets.
We evaluate on five multiple-choice QA benchmarks: OpenBookQA (OBQA) (Mihaylov et al., 2018), ARC-Easy (ARC-E) and ARC-Challenge (ARC-C) (Clark et al., 2018), BOOLQ (Clark et al., 2019), and ClimateQA—a domain-specific climate science benchmark. To assess robustness under distribution shift, we ensemble the domain-specific MMLU subtasks (Hendrycks et al., 2020) into $4$ benchmarks focusing on different professionalities: Computer Science (CS), Engineering (Eng), Law and Health. Details for the ensemble are provided in Appendix A.8. Models are fine-tuned on the public training split and evaluated on the test split for each benchmark.
Figure 3: Performance comparison of different methods fine-tuned with LlaMA2-7B across four common sense reasoning tasks and a domain-specific task. UQ4CT shows substantial ECE improvements while maintaining high accuracy.
Experiment Setup.
We implement UQ4CT with PyTorch (Paszke et al., 2019), extending the MixLoRA repository in (Li et al., 2024). We use the LLaMA-2-7B-hf (Touvron et al., 2023a) as our base model. In particular, we apply MixLoRA to query, key, value and output layers, together with the feed-forward networks in LLaMA-2-7B-hf (gate layer, down layer and up layer). Details are provided in Appendix A.6.
@@
  • Monte Carlo (MC) Dropout (Gal and Ghahramani, 2016) keeps dropout on at both training and testing time. By performing multiple forward passes, MC dropout randomly shuts down a portion of model nodes, producing ensemble-alike predictions. To combine LoRA fine-tuning with MC dropout, we apply dropout on the input of the LoRA adapter, following the implementation of Mangrulkar et al. (2022).
@@
  • Laplace-LoRA (LA) (Yang et al., 2024a) applies a post-hoc Laplace approximation on fine-tuned LoRA parameters for better uncertainty estimation.
@@
Evaluation.
We evaluate prediction accuracy on the validation set across all five tasks. For uncertainty calibration, we use the expected calibration error (ECE; Guo et al. (2017); more details in A.7) to measure the alignment between predicted probabilities and actual outcomes.
To assess robustness under distribution shifts, we fine-tune models on the OBQA dataset and evaluate them following (Yang et al., 2024a). We use ARC-C and ARC-E to represent smaller distribution shifts, as these datasets focus on general science reasoning similar to OBQA but are more challenging and diverse. For larger shifts, we utilize the four aforementioned domain-specific MMLU subtasks, which span a wide range of expertise from elementary to professional levels. This domain specificity represents a greater distribution shift from the general common sense focus of OBQA.
The in-distribution scenario tests model alignment on the target task, while the distribution shift scenario assesses generalizability to novel domains. Together, they provide a comprehensive evaluation for real-world applications, ensuring strong performance on the primary task and resilience to out-of-distribution inputs.
@@
MixLoRA $75.5_{2.91}$ $58.5_{1.44}$ $\mathbf{69.2_{1.02}}$ $35.2_{2.92}$ $30.3_{0.98}$ $\mathbf{35.9_{0.43}}$ $\mathbf{40.6_{1.13}}$
UQ4CT $\mathbf{77.3_{1.36}}$ $\mathbf{58.8_{1.06}}$ $65.8_{1.31}$ $\underline{36.2_{1.24}}$ $\mathbf{34.1_{2.31}}$ $\underline{35.8_{1.01}}$ $\underline{40.0_{1.24}}$
@@
5.1 In-distribution Performance
As shown in Figure 3, UQ4CT achieves notable gains in uncertainty calibration across diverse tasks, while maintaining competitive accuracy (ACC) relative to baseline approaches. For instance, on BoolQ and ClimateQA tasks, UQ4CT attains accuracy rates of $73.5\%$ and $63.3\%$, demonstrating that improved uncertainty quantification does not come at the expense of predictive performance.
The most significant improvements are seen in reduced Expected Calibration Error (ECE). UQ4CT consistently lowers ECE by over $25\%$ on average across benchmarks, and unlike other approaches, it continues to perform well even on challenging datasets such as ARC-C, achieving an ECE of $6.1$.
To further validate our method, we include results from fine-tuning both LLaMA-2-7B (main text) and Mistral-7B (Appendix A.3). Across both models, UQ4CT delivers substantial and consistent improvements in uncertainty calibration. These results underscore the practical value of UQ4CT, particularly in scenarios where reliable uncertainty estimates are crucial, such as safety-critical applications. A key advantage of UQ4CT is that it incorporates uncertainty calibration directly during fine-tuning, incurring minimal computational overhead compared to other uncertainty quantification (UQ) methods, which often require costly repetitive sampling or post-hoc adjustments.
5.2 Performance Under Distribution Shift
Due to the sparse nature of the fine-tuning data, real world deployment of LLMs often requires the model to be robust to out-of-distribution knowledge (Ouyang et al., 2022; Touvron et al., 2023d, a). Therefore, we evaluate the performance of UQ4CT along with other baseline models fine-tuned on the OBQA dataset under smaller and larger distribution shift scenarios.